### v1

@@ 3. DiCE-Extended CFs (Optimized CF Evaluation):
Figure 2: Loss trends in DiCE-Extended across datasets, illustrating the optimization of class loss, proximity loss, diversity loss, and robustness loss during CF generation.
Table 2: Classification performance of ML models trained and evaluated on original datasets.
Dataset Backend Accuracy F1-score Recall Precision AUC
Adult-income Scikit-learn 0.82 0.56 0.48 0.68 0.70
PYT 0.83 0.77 0.76 0.78 0.89
TF2 0.83 0.61 0.55 0.69 0.74
Compas-recidivism Scikit-learn 0.57 0.58 0.60 0.57 0.57
PYT 0.66 0.66 0.66 0.66 0.72
TF2 0.66 0.65 0.64 0.66 0.66
German-credit Scikit-learn 0.72 0.41 0.33 0.54 0.61
PYT 0.79 0.73 0.72 0.76 0.81
TF2 0.76 0.53 0.45 0.64 0.67
Lending-club Scikit-learn 0.82 0.90 0.99 0.83 0.51
PYT 0.83 0.46 0.50 0.60 0.69
TF2 0.83 0.91 1.00 0.83 0.51
Table 3: Classification performance of ML models trained on original datasets and evaluated on CF datasets (DiCE).
Dataset Backend Accuracy F1-score Recall Precision AUC Time (s) Validity
Adult-income Scikit-learn 1.00 1.00 1.00 1.00 1.00 1.00 0.77
PYT 0.56 0.53 0.56 0.58 0.55 0.59 0.89
TF2 0.41 0.11 0.07 0.22 0.41 0.52 0.88
Compas-recidivism Scikit-learn 1.00 1.00 1.00 1.00 1.00 1.00 0.49
PYT 1.00 1.00 1.00 1.00 1.00 0.88 0.18
TF2 0.44 0.41 0.38 0.44 0.44 0.68 0.49
German-credit Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.14 0.70
PYT 0.58 0.56 0.58 0.61 0.63 0.87 0.49
TF2 0.54 0.43 0.35 0.56 0.54 0.17 0.58
Lending-club Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.67 0.67
PYT 0.49 0.33 0.25 0.50 0.50 0.18 0.48
TF2 0.72 0.77 0.96 0.65 0.72 0.85 0.58
Table 4: Classification performance of ML models trained on original datasets and evaluated on CF datasets (DiCE-Extended).
Dataset Backend Accuracy F1-score Recall Precision AUC Time (s) Validity
Adult-income Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.20 0.89
PYT 0.47 0.39 0.47 0.43 0.46 0.61 0.97
TF2 0.51 0.29 0.20 0.54 0.52 0.27 0.99
Compas-recidivism Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.11 0.66
PYT 0.85 0.85 0.86 0.87 0.97 0.08 0.58
TF2 0.38 0.43 0.46 0.40 0.38 0.10 0.90
German-credit Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.49 0.45
PYT 0.61 0.61 0.62 0.63 0.69 0.10 0.48
TF2 0.55 0.42 0.33 0.59 0.55 0.64 0.99
Lending-club Scikit-learn 1.00 1.00 1.00 1.00 1.00 0.26 0.77
PYT 0.48 0.32 0.48 0.24 0.23 0.18 0.76
TF2 0.70 0.76 0.91 0.65 0.71 0.10 0.87
@@ 5 Conclusion
Our quantitative analysis highlights that DiCE-Extended improves the robustness, validity, and overall reliability of CF explanations across different ML backends. While DiCE-generated CFs introduce notable performance degradation, DiCE-Extended mitigates these issues, ensuring better alignment with decision boundaries and producing more actionable CFs. Future work will focus on optimizing multi-objective trade-offs, integrating domain-specific constraints, and improving computational efficiency, making DiCE-Extended a more practical and interpretable counterfactual generation framework for real-world applications.

### v2

@@ 3.1 Loss Progression in CF Explainers
Figure 1: Loss trends in DiCE-Extended across datasets for the TensorFlow backend, illustrating the optimization of class loss, proximity loss, diversity loss, and robustness loss during CF generation.
The convergence of robustness loss is aligned with the overall minimization strategy, which simultaneously balances proximity, diversity, and stability without compromising interpretability. Furthermore, the observed stabilization suggests that after a certain number of iterations, further optimization yields diminishing returns in terms of robustness improvement. This finding highlights the efficiency of DiCE-Extended in generating robust CFs within a limited number of optimization steps, making it suitable for time-sensitive and decision-critical applications.
3.2 DiCE-Extended vs. DiCE: Counterfactual Quality and Fidelity
To assess the added value of DiCE-Extended, we evaluated several performance metrics, grouped into the following categories:
1. Counterfactual Quality Suite. For each method, we report standard evaluation metrics, including validity, proximity, sparsity, diversity, and the proposed robustness measure, as illustrated in Figure 1.
2. Local Fidelity. To evaluate local fidelity, we trained a 1-nearest-neighbour (1-NN) surrogate model on the union of the original query point and its generated CFs. The surrogate’s accuracy was then measured on 1000 synthetic neighbors sampled and scaled using the mean absolute deviation (MAD) around the query. The MAD is defined as:
$\text{MAD}=\frac{1}{n}\sum_{i=1}^{n}\lvert x_{i}-m(X)\rvert,$
where $m(X)$ represents the mean of the dataset $X$. Comparative local fidelity results for DiCE-Extended and the benchmark DiCE method are reported in Table 2 and Table 3. Higher 1-NN accuracy indicates that the CF set more accurately reflects the model’s local decision boundary [19].
Counterfactual Generation.
For each (dataset, backend) pair, we randomly sampled $N=50$ test instances and generated $k=10$ CFs for each instance, targeting the opposite class (desired_class $=1-y$).
Statistical Analysis.
For each instance, metric differences between DiCE-Extended and DiCE
[…]
were analyzed using two-tailed paired $t$-tests ($H_{0}$: mean difference $=0$; $\alpha=0.05$). $p$-values are reported in scientific notation, and values below 0.05 are highlighted in bold to indicate statistical significance.
3.3 Performance Gains Across Counterfactual Dimensions
DiCE-Extended consistently generated closer yet valid CFs, as shown in Table 2. Across all four datasets and three model back-ends, the median proximity improved significantly, shrinking by 30%–95% while maintaining a validity rate of $100\%$. A two-tailed paired $t$-test on the 50 per-instance proximity scores confirmed that these reductions were highly significant ($p<10^{-5}$).
The framework also achieved notable improvements in diversity and sparsity. On high-dimensional datasets such as Adult and Lending-Club, CF diversity nearly doubled. Concurrently, Sparsity, defined as the average number of features altered, showed improvements of up to $50\%$, with all differences statistically significant at the p-value threshold of $<0.01$.
Improvements were most pronounced in the newly introduced robustness metric. For PyTorch and TensorFlow models, DiCE-Extended demonstrated absolute gains of 0.05–0.13 in robustness scores, all statistically significant at the $p<0.01$ level. These results suggest that the additional robustness loss effectively safeguards CF explanations against input perturbations. Although the gains were smaller for tree-based Scikit-learn models, which are inherently more stable, they were directionally consistent.
Importantly, these enhancements were achieved with only minimal computational overhead due to vectorized implementation. On average, generation time increased by approximately 1.2×, yet each configuration successfully produced the required number of valid CFs per query without degradation in performance.
The 1-NN fidelity evaluation further demonstrated that DiCE-Extended preserves, and in many cases improves, the alignment of CFs with the model’s local decision boundary. Specifically, 1-NN fidelity increased by 5–11% for Scikit-learn models on the Adult-Income and Lending-Club datasets, with statistical significance at the 2% level. For TensorFlow on German-Credit, improvements of 3–4% were similarly significant. In contrast, PyTorch on COMPAS showed a modest 7% advantage for the original DiCE method, but this difference was not statistically significant. In all other configurations, differences were below 4% and statistically indistinguishable from zero.
Table 2: 1-NN fidelity scores (mean $\pm$ standard deviation, $N=50$) for counterfactual sets generated by DiCE-Extended across three MAD radii (0.5, 1.0, and 2.0).
Dataset Backend 0.5 MAD 1.0 MAD 2.0 MAD
Adult-Income sklearn 0.54 $\pm$ 0.20 0.48 $\pm$ 0.18 0.43 $\pm$ 0.16
PYT 0.54 $\pm$ 0.16 0.53 $\pm$ 0.16 0.51 $\pm$ 0.17
TF2 0.54 $\pm$ 0.069 0.54 $\pm$ 0.078 0.54 $\pm$ 0.076
COMPAS-Recidivism sklearn 0.53 $\pm$ 0.12 0.53 $\pm$ 0.11 0.52 $\pm$ 0.087
PYT 0.57 $\pm$ 0.20 0.56 $\pm$ 0.19 0.55 $\pm$ 0.16
TF2 0.89 $\pm$ 0.076 0.89 $\pm$ 0.074 0.88 $\pm$ 0.075
German-Credit sklearn 0.49 $\pm$ 0.20 0.42 $\pm$ 0.20 0.38 $\pm$ 0.19
PYT 0.48 $\pm$ 0.13 0.48 $\pm$ 0.13 0.46 $\pm$ 0.13
TF2 0.58 $\pm$ 0.052 0.58 $\pm$ 0.047 0.59 $\pm$ 0.046
Lending-Club sklearn 0.33 $\pm$ 0.16 0.26 $\pm$ 0.14 0.22 $\pm$ 0.11
PYT 0.46 $\pm$ 0.08 0.44 $\pm$ 0.081 0.42 $\pm$ 0.078
TF2 0.49 $\pm$ 0.056 0.47 $\pm$ 0.055 0.43 $\pm$ 0.067
Note: Highest fidelity per dataset at each MAD radius is highlighted in bold.
Table 3: 1-NN fidelity scores (mean $\pm$ standard deviation, $N=50$) for counterfactual sets generated by DiCE across three MAD radii (0.5, 1.0, and 2.0).
Dataset Backend 0.5 MAD 1.0 MAD 2.0 MAD
Adult-Income sklearn 0.43 $\pm$ 0.18 0.40 $\pm$ 0.20 0.38 $\pm$ 0.21
PYT 0.58 $\pm$ 0.18 0.57 $\pm$ 0.18 0.56 $\pm$ 0.19
TF2 0.54 $\pm$ 0.085 0.53 $\pm$ 0.080 0.51 $\pm$ 0.086
COMPAS-Recidivism sklearn 0.62 $\pm$ 0.14 0.56 $\pm$ 0.14 0.55 $\pm$ 0.15
PYT 0.64 $\pm$ 0.23 0.62 $\pm$ 0.21 0.61 $\pm$ 0.18
TF2 0.92 $\pm$ 0.081 0.91 $\pm$ 0.085 0.88 $\pm$ 0.10
German-Credit sklearn 0.44 $\pm$ 0.20 0.38 $\pm$ 0.21 0.35 $\pm$ 0.22
PYT 0.45 $\pm$ 0.15 0.44 $\pm$ 0.15 0.43 $\pm$ 0.15
TF2 0.55 $\pm$ 0.068 0.55 $\pm$ 0.062 0.55 $\pm$ 0.051
Lending-Club sklearn 0.24 $\pm$ 0.11 0.20 $\pm$ 0.11 0.18 $\pm$ 0.11
PYT 0.52 $\pm$ 0.084 0.51 $\pm$ 0.076 0.48 $\pm$ 0.078
TF2 0.47 $\pm$ 0.092 0.44 $\pm$ 0.100 0.40 $\pm$ 0.100
Note: Highest fidelity per dataset at each MAD radius is highlighted in bold.
@@ 4 Conclusion
Our findings demonstrate that, particularly for DL back-ends, DiCE-Extended consistently generates CFs that are closer, sparser, and more diverse, while maintaining near-perfect validity and significantly enhancing robustness. Robustness scores increased by $+0.05$ to $+0.13$ (with $p<0.05$), and substantial gains in proximity indicate that the generated CFs more faithfully adhere to the model’s local decision boundaries. Despite a modest increase in generation time (approximately 1.2×1.2$\times$1.2 ×), each configuration successfully produced the desired number of CFs per query. Future research will focus on fine-tuning the trade-offs among proximity, diversity, and robustness; integrating domain-specific constraints; and extending the framework with robust data privacy guarantees, thereby enabling privacy-aware CF generation in real-world applications.